# Efficient GW band structure calculations using Gaussian basis functions and application to atomically thin transition-metal dichalcogenides

### v2

@@ I Introduction
GW calculations [1, 2, 3] have become a standard method for calculating electron addition and removal energies of molecules [4, 5, 6, 7], two-dimensional materials [8, 9, 10] and bulk solids [11, 12, 13]. Recent advancements of the $GW$ method span a broad spectrum, including the application to deep core excitations [14, 15, 16, 17, 18, 19, 20, 21, 22], relativistic GW schemes [23, 24, 25, 26, 27, 28] and vertex corrections [29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. These developments have firmly established GW as a powerful and versatile approach within the domain of many-body perturbation theory.
@@
Despite the methodological maturity of $GW$, computational challenges remain, particularly when applied to atomically thin materials. One of the primary limitations arises from the use of plane-wave basis sets in systems with vacuum, such as molecules or low-dimensional systems. The need to represent the vacuum leads to large plane-wave basis set size and significant computational cost. These constraints motivate the development of alternative basis representations. One compelling solution involves atom-centered basis functions, which are localized and naturally adapted to such geometries. Atom-centered basis functions are already the standard for $GW$ implementations targeting molecules [4, 6, 7, 3, 5]. Also, several implementations of $GW$ with atom-centered basis functions and periodic boundary conditions have been reported [39, 40, 41, 42, 43, 44, 45, 46, 47]. The first pioneering $GW$ calculations in semiconductors were carried by Strinati et al. [39, 40], who studied the band structure of diamond using a local-orbital formulation of the $GW$ calculation. In the 90ies, Rohlfing et al. [41, 42] studied bulk semiconductors and a silicon surface finding good agreement to plane-wave based $GW$ implementations in the order of 0.1 eV or better. The more recent periodic $GW$ implementations with atom-centered basis functions [43, 44, 45, 46] focus on 3D crystals and report similar precision. The $GW$ implementations [39, 40, 41, 42, 43, 44, 45, 46] use a formulation of $GW$ in frequency and rely on $k$-point sampling to account for the periodic boundary conditions.
@@
An alternative is the $GW$ space-time method [48, 49] which has been originally formulated using plane waves and real-space representations as well as time and frequency representations. The $GW$ space-time method can be also reformulated using an atom-centered basis set [50, 51, 52, 53] which eliminates the use of plane-waves and real-space grids and allows for low-scaling $GW$ calculations on large molecules. The $GW$ space-time approach in an atom-centered basis can be also combined with periodic boundary conditions, as we have demonstrated in our previous work [47] where we employed a $\Gamma$-only approach for the density response function and the self-energy, while relying on dense $k$-point sampling for the screened Coulomb interaction $W$ to account for the divergence of $W$ at the $\Gamma$-point. This implementation enabled the study of a twisted transition-metal dichalcogenide heterobilayers with almost 1000 atoms in the unit cell. The drawback of the $\Gamma$-only implementation [47] is that large unit cells are required where the density response needs to vanish on the length scale of the unit cell.
@@
In this work, we address the limitation of the $\Gamma$-point only approach [47], which requires large unit cells, by introducing a lattice summation over neighbor cells for both the density response and self-energy, as originally proposed in the $GW$ space-time method [48]. This extension enables the accurate and efficient treatment of crystals with small unit cells. The resulting $GW$ algorithm is particularly well-suited for low-dimensional materials, as the number of neighboring cells required in the lattice sums is significantly reduced compared to 3D bulk crystals. Furthermore, the Gaussian basis set is efficient in simulations involving large vacuum regions, a common requirement when modeling low-dimensional systems.
@@
Our implementation also supports the inclusion of relativistic effects via spin-orbit coupling (SOC) from Gaussian dual-space pseudopotentials [54, 55, 56] and a perturbative correction to the quasiparticle energies. We focus on $GW$ band structure calculations of atomically thin transition metal dichalcogenides (TMDCs) MoS2, MoSe2, WS2, WSe2 and we benchmark our $GW$ band structures against state-of-the-art plane-wave-based $GW$ implementations in BerkeleyGW [57] and VASP [12]. We demonstrate that our $GW$ algorithm yields accurate and converged quasiparticle band structures across multiple convergence parameters, including basis set size, $k$-point sampling, the summation of neighbor cells, and the time- and frequency-mesh. We also discuss the computational efficiency of our $GW$ band structure algorithm, which enables a $GW$ band structure calculation of an atomically thin material on a laptop (Intel Xeon i5, 192 GB RAM) within roughly a day.
@@ II GW space-time algorithm
Many efficient $GW$ algorithms [49, 51, 52] build on the $GW$ space-time method [48]. In order to introduce the basic idea of the $GW$ space-time method, we use a generic formulation in this section for non-periodic systems projecting all quantities on real-space grids. It is important to note that this formulation differs from the original $GW$ space-time method [48] where some quantities were calculated using a plane-wave basis set.
In this work, we employ the $G_{0}W_{0}$ scheme which starts from a self-consistent Kohn-Sham density functional theory (KS-DFT) calculation [58],
@@
This transform can be understood as Laplace transform followed by analytic continuation to the imaginary axis and effectively takes the form of a Fourier transform [59]. Next, the dielectric function $\epsilon$ can be calculated in imaginary frequency from the irreducible polarizability,
@@
The self-energy is then analytically continued to real frequency [3, 60].
@@ III KS-DFT with periodic boundary conditions and Gaussian basis functions
We use KS-DFT with periodic boundary conditions [61, 62, 63], i.e., $h_{0}(\mathbf{r})$ and $v^{\text{xc}}(\mathbf{r})$ from Eq. (1) are lattice periodic,
@@
Bloch’s theorem [64] states that the solutions $\psi_{n\mathbf{k}}(\mathbf{r})$ of the Kohn-Sham equations
@@
The integration over the BZ in Eq. (23) is executed in every self-consistent field cycle of the KS-DFT calculation using a discrete $N_{1}\hskip 0.89996pt{\times}\hskip 0.59998ptN_{2}\hskip 0.89996pt{\times}\hskip 0.59998ptN_{3}$ Monkhorst-Pack $k$-point mesh $\{\mathbf{k}_{\ell}\}$ [65] for two-dimensional periodicity:
@@
Starting from the eigenvalues $\varepsilon_{n\mathbf{k}}^{\operatorname{DFT}}$ from Eq. (18), we add spin-orbit coupling (SOC) $V^{\operatorname{SOC}}$ from Hartwigsen-Goedecker-Hutter (HGH) pseudopotentials [54, 55, 56], see details in Appendix H, to obtain the spin-orbit perturbed Hamiltonian:
@@ IV GW space-time method with Gaussian basis functions and periodic boundary conditions
As shown in Appendix B, the Green’s function in imaginary time is given by [49]
@@
For the computation of the density response $\chi$, we introduce the Resolution of Identity (RI) [66], where products of Gaussian basis functions are expanded over an auxiliary RI basis set $\{\varphi_{P}^{\mathbf{P}}\}$:
@@
that can be understood as projection of the product $\phi_{\mu}^{\mathbf{R}_{1}}\phi_{\nu}^{\mathbf{R}_{2}}$ onto the RI basis function $\varphi_{P}^{\mathbf{R}}$ under a metric, which we have chosen to be the truncated Coulomb operator (RI-tCm) [67, 68, 69],
@@
The cutoff radius $r_{\text{c}}$ from Eq. (40) is typically chosen in the order of a few Ångstroms [68, 69, 47]. In the limit $r_{\text{c}}\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt0$, RI-tCm is equivalent to the RI with the overlap metric which suffers from a slow convergence with respect to the RI basis set size [66, 67]. In the limit of $r_{\text{c}}\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$, RI-tCM is equivalent to the RI with the Coulomb metric where the convergence with the RI basis set size is fast. The drawback of the Coulomb metric is that tensor elements $(\mu\mathbf{R}\,\nu\mathbf{T}\,|\,P\mathbf{P})$ from Eq. (37) with the Coulomb operator only decay polynomially for $|\mathbf{R}{-}\mathbf{P}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$ and $|\mathbf{T}{-}\mathbf{P}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$ which prohibits the application of RI with the Coulomb metric in this algorithm. When truncating the Coulomb operator as in Eq. (40) at finite $r_{\text{c}}$ and when using Gaussian basis functions, the tensor $(\mu\mathbf{R}\,\nu\mathbf{T}\,|\,P\mathbf{P})$ decays like a Gaussian for $|\mathbf{R}{-}\mathbf{P}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$ and $|\mathbf{T}{-}\mathbf{P}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$. It has been shown that RI-tCm converges quickly with the size of the auxiliary basis $\{\varphi_{P}^{\mathbf{R}}\}$[67, 68]
@@
Following the $GW$ space-time method, [48, 47] we transform the polarizability from real space and time to the Brillouin zone and frequency
@@
Note that as $\chi^{\mathbf{R}}$ computed from Eq. (42) decays exponentially for large $\mathbf{R}$ in gapped systems, so one can perform the Fourier transformation (43) to an arbitrary $\mathbf{k}$. We execute the imaginary-time integration numerically using minimax grids [49, 70, 71].
As next step, we calculate the dielectric function [47]
@@
where $\mathbf{Id}$ is the identity matrix and the truncated Coulomb matrix $\mathbf{M}(\mathbf{k})$ defined in Eq. (41) appears due to the RI-tCm. We use Tikhonov regularization [72] for the RI expansion to prevent linear dependencies of fit coefficients, as discussed in the supporting information of Ref. [47]. The regularization leads to the modified matrix inversion,
@@
$\mathbf{V}^{0.5}(\mathbf{k})$ in Eq. (44) is the matrix square root of the bare Coulomb interaction $\mathbf{V}(\mathbf{k})$ [73, 44, 74, 75]
@@
Special care is required for the BZ integral as $W_{PQ}^{\text{c}}(\mathbf{k},i\omega)$ diverges at the $\Gamma$-point with $1/k$ for two-dimensional materials if $\varphi_{P}$ and $\varphi_{Q}$ are s-type basis functions [43, 45, 44]. We evaluate $W_{PQ}^{\text{c}}(\mathbf{k},i\omega)$ using two Monkhorst-Pack meshes [65]: $\{\mathbf{k}_{\ell}\}$ has $4N_{j}$ $k$-points in periodic directions $j$ and $\{\mathbf{q}_{\ell}\}$ has $8N_{j}$ $k$-points in periodic directions $j$. The number of $k$-points in the $\{\mathbf{k}_{\ell}\}$ mesh and $\{\mathbf{q}_{\ell}\}$ mesh is thus
@@
We extrapolate the BZ integration (48) with the inverse square root of the number of $k$-points [44]. Reformulating Eq. (48) , the $k$-extrapolated screened Coulomb interaction becomes
@@
${\color[rgb]{0,0,0}\simeq\frac{1}{N_{1}N_{2}N_{3}}\sum_{\mathbf{k}}^{\textrm{BZ}}e^{-i\mathbf{k}_{\ell}\cdot\mathbf{R}}\,\tilde{W}^{\text{c}}_{PQ}(\mathbf{k}_{\ell},i\omega)\,.}$ (55)
Following the $GW$ space-time method [48], we transform $W^{\text{c}}(i\omega)$ to imaginary time using minimax grids [49, 70, 71]. This completes the ingredients for the $GW$ self-energy $\Sigma^{\text{c}}(i\tau)\hskip 1.69998pt{=}\hskip 1.69998ptiG(i\tau)W^{\text{c}}(i\tau)$. We calculate the self-energy $\Sigma^{\text{c},\mathbf{R}}_{\lambda\sigma}(i\tau)\hskip 1.69998pt{=}\hskip 1.69998pt\braket{\phi_{\lambda}^{\mathbf{0}}|\Sigma^{\text{c}}(i\tau)|\phi_{\sigma}^{\mathbf{R}}}$ in real space (derivation in Appendix G),
@@
A major advantage of the present $GW$ algorithm is that the self-energy is computed in real space, Eq. (56), involving only a modest number of neighbor cells $\mathbf{R}_{1},\mathbf{R}_{2},\mathbf{S}_{1},\mathbf{S}_{2}$ to be included. Instead, when using $k$-point sampling instead of lattice summation to compute the self-energy, special care needs to be taken due to the diverging $W(\mathbf{k})$ at the $\Gamma$-point, which requires special correction schemes [76]. In the present algorithm, the divergence of $W(\mathbf{k})$ at the $\Gamma$-point is taken into account via BZ integration of $W(\mathbf{k})$ in Eq. (48) with suitable $k$-point grids, Eq. (50).
We calculate the exchange self-energy similarly to previous work [77, 69],
$\Sigma^{\text{x},\mathbf{R}}_{\lambda\sigma}(i\tau)$ $=i\sum_{P\nu}\sum_{\mathbf{R}_{1}}^{\text{SC}}\sum_{\mathbf{S}_{1}}^{\text{SC}}\left[\sum_{\mu}\sum_{\mathbf{S}_{2}}^{\text{SC}}(\lambda\mathbf{0}\;\mu\mathbf{S}_{1}{-}\mathbf{S}_{2}\,|\,P\mathbf{R}_{1})\;D^{\mathbf{S}_{2}}_{\mu\nu}\right]$
@@
is the truncated Coulomb matrix with truncation radius $r_{\text{SC}}\hskip 1.69998pt{=}\hskip 1.69998pt\min_{j}N_{j}|\mathbf{a}_{j}|/2$, i.e., half of the shortest primitive translation vector of the supercell SC (29[78, 77]. Note that we restrict all lattice vector differences appearing in Eq. (56) and (59) to the SC (29), i.e., $\mathbf{S}_{1}{-}\mathbf{S}_{2},\mathbf{R}_{1}{-}\mathbf{R}_{2}\hskip 1.69998pt{\in}\hskip 1.69998pt\text{SC}$.
@@
The primary difference between our algorithm and the $\Gamma$-point-only $GW$ implementation using Gaussian basis functions reported in Ref. [47] lies in the treatment of periodic boundary conditions. In our approach, both the irreducible density response function $\chi$ and the self-energy $\Sigma$ are computed via explicit lattice summations [Eqs. (42) and (56)] over all unit cells. In contrast, Ref. [47] restricts the lattice sum to the nearest-neighbor cell, effectively circumventing the need for a full lattice summation. This approach is only exact in the limit of large unit cells, and the presented $GW$ algorithm enables the treatment of small unit cells.

### v3

@@ I Introduction
GW calculations [hedin1965new, Reining2018, Golze2019] have become a standard method for calculating electron addition and removal energies of molecules [Blase2011, Faber2011, gw100, Knight2016], two-dimensional materials [Molina2013, Gjerding2021, Altman2024] and bulk solids [bruneval2008accurate, Klimes2014, Leppert2024]. Recent advancements of the $GW$ method span a broad spectrum, including the application to deep core excitations [Golze2018, Golze2020, keller2020relativistic, mejia2021, Mejia2022basis, li2022benchmark, panades2023, Galleni2024, Kocklaeuner2025], relativistic GW schemes [holzer2019ionized, Yeh2022, foerster2023twocomponent, kehry2023, Gaurav2024, Abraham2024] and vertex corrections [Grueneis2014, Maggio2017Vertex, Lewis2019, Tal2021, wangy2021, Foerster2022ExploringG3W2, Lorin2023, bruneval2024fully, Wen2024, Foerster2025]. These developments have firmly established GW as a powerful and versatile approach within the domain of many-body perturbation theory.
@@
Despite the methodological maturity of $GW$, computational challenges remain, particularly when applied to atomically thin materials. One of the primary limitations arises from the use of plane-wave basis sets in systems with vacuum, such as molecules or low-dimensional systems. The need to represent the vacuum leads to large plane-wave basis set size and significant computational cost. These constraints motivate the development of alternative basis representations. One compelling solution involves atom-centered basis functions, which are localized and naturally adapted to such geometries. Atom-centered basis functions are already the standard for $GW$ implementations targeting molecules [Blase2011, gw100, Knight2016, Golze2019, Faber2011]. Also, several implementations of $GW$ with atom-centered basis functions and periodic boundary conditions have been reported [Strinati1980, Strinati1982, Rohlfing1993, Rohlfing1995, Wilhelm2017, zhu2021all, Ren/etal:2021, Harsha2024, Graml2024]. The first pioneering $GW$ calculations in semiconductors were carried out by Strinati et al. [Strinati1980, Strinati1982], who studied the band structure of diamond using a local-orbital formulation of the $GW$ calculation. In the 90ies, Rohlfing et al. [Rohlfing1993, Rohlfing1995] studied bulk semiconductors and a silicon surface finding good agreement to plane-wave based $GW$ implementations in the order of 0.1 eV or better. The more recent periodic $GW$ implementations with atom-centered basis functions [Wilhelm2017, zhu2021all, Ren/etal:2021, Harsha2024] focus on 3D crystals and report similar precision. The $GW$ implementations [Strinati1980, Strinati1982, Rohlfing1993, Rohlfing1995, Wilhelm2017, zhu2021all, Ren/etal:2021, Harsha2024] use a formulation of $GW$ in frequency and rely on $k$-point sampling to account for the periodic boundary conditions.
@@
An alternative is the $GW$ space-time method [rojas_1995, liu2016] which has been originally formulated using plane waves and real-space representations as well as time and frequency representations. The $GW$ space-time method can be also reformulated using an atom-centered basis set [wilhelm2018, foerster2020, duchemin2021, foerster2022quasiparticle] which eliminates the use of plane-waves and real-space grids and allows for low-scaling $GW$ calculations on large molecules. The $GW$ space-time approach in an atom-centered basis can be also combined with periodic boundary conditions, as we have demonstrated in our previous work [Graml2024] where we employed a $\Gamma$-only approach for the density response function and the self-energy, while relying on dense $k$-point sampling for the screened Coulomb interaction $W$ to account for the divergence of $W$ at the $\Gamma$-point. This implementation enabled the study of a twisted transition-metal dichalcogenide heterobilayers with almost 1000 atoms in the unit cell. The drawback of the $\Gamma$-only implementation [Graml2024] is that large unit cells are required where the density response needs to vanish on the length scale of the unit cell.
@@
In this work, we address the limitation of the $\Gamma$-point only approach [Graml2024], which requires large unit cells, by introducing a lattice summation over neighbor cells for both the density response and self-energy, as originally proposed in the $GW$ space-time method [rojas_1995]. This extension enables the accurate and efficient treatment of crystals with small unit cells. The resulting $GW$ algorithm is particularly well-suited for low-dimensional materials, as the number of neighboring cells required in the lattice sums is significantly reduced compared to 3D bulk crystals. Furthermore, the Gaussian basis set is efficient in simulations involving large vacuum regions, a common requirement when modeling low-dimensional systems.
@@
Our implementation also supports the inclusion of relativistic effects via spin-orbit coupling (SOC) from Gaussian dual-space pseudopotentials [Hartwigsen1998, Krack2005, Vogt2025] and a perturbative correction to the quasiparticle energies. We focus on $GW$ band structure calculations of atomically thin transition metal dichalcogenides (TMDCs) MoS2, MoSe2, WS2, WSe2 and we benchmark our $GW$ band structures against state-of-the-art plane-wave-based $GW$ implementations in BerkeleyGW [Deslippe2012] and VASP [Klimes2014]. We demonstrate that our $GW$ algorithm yields accurate and converged quasiparticle band structures across multiple convergence parameters, including basis set size, $k$-point sampling, the summation of neighbor cells, and the time- and frequency-mesh. We also discuss the computational efficiency of our $GW$ band structure algorithm, which enables a $GW$ band structure calculation of an atomically thin material on a laptop (Intel i5 13600HX, 192 GB RAM) within roughly a day.
@@ II GW space-time algorithm
Many efficient $GW$ algorithms [liu2016, foerster2020, duchemin2021] build on the $GW$ space-time method [rojas_1995]. In order to introduce the basic idea of the $GW$ space-time method, we use a generic formulation in this section for non-periodic systems projecting all quantities on real-space grids. It is important to note that this formulation differs from the original $GW$ space-time method [rojas_1995] where some quantities were calculated using a plane-wave basis set.
In this work, we employ the $G_{0}W_{0}$ scheme which starts from a self-consistent Kohn-Sham density functional theory (KS-DFT) calculation [Kohn1965],
@@
This transform can be understood as Laplace transform followed by analytic continuation to the imaginary axis and effectively takes the form of a Fourier transform [rieger1999gw]. Next, the dielectric function $\epsilon$ can be calculated in imaginary frequency from the irreducible polarizability,
@@
The self-energy is then analytically continued to real frequency [Golze2019, Leucke2025].
@@ III KS-DFT with periodic boundary conditions and Gaussian basis functions
We use KS-DFT with periodic boundary conditions [Delley2000, Martin2004, Blum2009], i.e., $h_{0}(\mathbf{r})$ and $v^{\text{xc}}(\mathbf{r})$ from Eq. (1) are lattice periodic,
@@
Bloch’s theorem [Bloch1929] states that the solutions $\psi_{n\mathbf{k}}(\mathbf{r})$ of the Kohn-Sham equations
@@
The integration over the BZ in Eq. (23) is executed in every self-consistent field cycle of the KS-DFT calculation using a discrete $N_{1}\hskip 0.89996pt{\times}\hskip 0.59998ptN_{2}\hskip 0.89996pt{\times}\hskip 0.59998ptN_{3}$ Monkhorst-Pack $k$-point mesh $\{\mathbf{k}_{\ell}\}$ [Monkhorst1976] for two-dimensional periodicity:
@@
Starting from the eigenvalues $\varepsilon_{n\mathbf{k}}^{\operatorname{DFT}}$ from Eq. (18), we add spin-orbit coupling (SOC) $V^{\operatorname{SOC}}$ from Hartwigsen-Goedecker-Hutter (HGH) pseudopotentials [Hartwigsen1998, Krack2005, Vogt2025], see details in Appendix H, to obtain the spin-orbit perturbed Hamiltonian:
@@ IV GW space-time method with Gaussian basis functions and periodic boundary conditions
As shown in Appendix B, the Green’s function in imaginary time is given by [liu2016]
@@
For the computation of the density response $\chi$, we introduce the Resolution of Identity (RI) [Vahtras1993], where products of Gaussian basis functions are expanded over an auxiliary RI basis set $\{\varphi_{P}^{\mathbf{P}}\}$:
@@
that can be understood as projection of the product $\phi_{\mu}^{\mathbf{R}_{1}}\phi_{\nu}^{\mathbf{R}_{2}}$ onto the RI basis function $\varphi_{P}^{\mathbf{R}}$ under a metric, which we have chosen to be the truncated Coulomb operator (RI-tCm) [Jung2005, wilhelm2021, Bussy2024],
@@
The cutoff radius $r_{\text{c}}$ from Eq. (40) is typically chosen in the order of a few Ångstroms [wilhelm2021, Bussy2024, Graml2024]. In the limit $r_{\text{c}}\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt0$, RI-tCm is equivalent to the RI with the overlap metric which suffers from a slow convergence with respect to the RI basis set size [Vahtras1993, Jung2005]. In the limit of $r_{\text{c}}\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$, RI-tCM is equivalent to the RI with the Coulomb metric where the convergence with the RI basis set size is fast. The drawback of the Coulomb metric is that tensor elements $(\mu\mathbf{R}\,\nu\mathbf{T}\,|\,P\mathbf{P})$ from Eq. (37) with the Coulomb operator only decay polynomially for $|\mathbf{R}{-}\mathbf{P}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$ and $|\mathbf{T}{-}\mathbf{P}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$ which prohibits the application of RI with the Coulomb metric in this algorithm. When truncating the Coulomb operator as in Eq. (40) at finite $r_{\text{c}}$ and when using Gaussian basis functions, the tensor $(\mu\mathbf{R}\,\nu\mathbf{T}\,|\,P\mathbf{P})$ decays like a Gaussian for $|\mathbf{R}{-}\mathbf{P}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$ and $|\mathbf{T}{-}\mathbf{P}|\hskip 1.00006pt{\rightarrow}\hskip 1.00006pt\infty$. It has been shown that RI-tCm converges quickly with the size of the auxiliary basis $\{\varphi_{P}^{\mathbf{R}}\}$[Jung2005, wilhelm2021]
@@
Following the $GW$ space-time method, [rojas_1995, Graml2024] we transform the polarizability from real space and time to the Brillouin zone and frequency
@@
Note that as $\chi^{\mathbf{R}}$ computed from Eq. (42) decays exponentially for large $\mathbf{R}$ in gapped systems, so one can perform the Fourier transformation (43) to an arbitrary $\mathbf{k}$. We execute the imaginary-time integration numerically using minimax grids [liu2016, Azizi2023, Azizi2024].
As next step, we calculate the dielectric function [Graml2024]
@@
where $\mathbf{Id}$ is the identity matrix and the truncated Coulomb matrix $\mathbf{M}(\mathbf{k})$ defined in Eq. (41) appears due to the RI-tCm. We use Tikhonov regularization [Goncharsky1995] for the RI expansion to prevent linear dependencies of fit coefficients, as discussed in the supporting information of Ref. [Graml2024]. The regularization leads to the modified matrix inversion,
@@
$\mathbf{V}^{0.5}(\mathbf{k})$ in Eq. (44) is the matrix square root of the bare Coulomb interaction $\mathbf{V}(\mathbf{k})$ [Grundei2017, zhu2021all, Garcia2025, spadetto2025]
@@
Special care is required for the BZ integral as $W_{PQ}^{\text{c}}(\mathbf{k},i\omega)$ diverges at the $\Gamma$-point with $1/k$ for two-dimensional materials if $\varphi_{P}$ and $\varphi_{Q}$ are s-type basis functions [Wilhelm2017, Ren/etal:2021, zhu2021all]. We evaluate $W_{PQ}^{\text{c}}(\mathbf{k},i\omega)$ using two Monkhorst-Pack meshes [Monkhorst1976]: $\{\mathbf{k}_{\ell}\}$ has $4N_{j}$ $k$-points in periodic directions $j$ and $\{\mathbf{q}_{\ell}\}$ has $8N_{j}$ $k$-points in periodic directions $j$. The number of $k$-points in the $\{\mathbf{k}_{\ell}\}$ mesh and $\{\mathbf{q}_{\ell}\}$ mesh is thus
@@
We extrapolate the BZ integration (48) with the inverse square root of the number of $k$-points [zhu2021all]. Reformulating Eq. (48) , the $k$-extrapolated screened Coulomb interaction becomes
@@
${\color[rgb]{0,0,0}\simeq\frac{1}{N_{1}N_{2}N_{3}}\sum_{\mathbf{k}_{\ell}}^{\textrm{BZ}}e^{-i\mathbf{k}_{\ell}\cdot\mathbf{R}}\,\tilde{W}^{\text{c}}_{PQ}(\mathbf{k}_{\ell},i\omega)\,.}$ (55)
Following the $GW$ space-time method [rojas_1995], we transform $W^{\text{c}}(i\omega)$ to imaginary time using minimax grids [liu2016, Azizi2023, Azizi2024]. This completes the ingredients for the $GW$ self-energy $\Sigma^{\text{c}}(i\tau)\hskip 1.69998pt{=}\hskip 1.69998ptiG(i\tau)W^{\text{c}}(i\tau)$. We calculate the self-energy $\Sigma^{\text{c},\mathbf{R}}_{\lambda\sigma}(i\tau)\hskip 1.69998pt{=}\hskip 1.69998pt\braket{\phi_{\lambda}^{\mathbf{0}}|\Sigma^{\text{c}}(i\tau)|\phi_{\sigma}^{\mathbf{R}}}$ in real space (derivation in Appendix G),
@@
A major advantage of the present $GW$ algorithm is that the self-energy is computed in real space, Eq. (56), involving only a modest number of neighbor cells $\mathbf{R}_{1},\mathbf{R}_{2},\mathbf{S}_{1},\mathbf{S}_{2}$ to be included. Instead, when using $k$-point sampling instead of lattice summation to compute the self-energy, special care needs to be taken due to the diverging $W(\mathbf{k})$ at the $\Gamma$-point, which requires special correction schemes [FREYSOLDT20071]. In the present algorithm, the divergence of $W(\mathbf{k})$ at the $\Gamma$-point is taken into account via BZ integration of $W(\mathbf{k})$ in Eq. (48) with suitable $k$-point grids, Eq. (50).
We calculate the exchange self-energy similarly to previous work [Irmler2018, Bussy2024],
$\Sigma^{\text{x},\mathbf{R}}_{\lambda\sigma}$ $=i\sum_{P\nu}\sum_{\mathbf{R}_{1}}^{\text{SC}}\sum_{\mathbf{S}_{1}}^{\text{SC}}\left[\sum_{\mu}\sum_{\mathbf{S}_{2}}^{\text{SC}}(\lambda\mathbf{0}\;\mu\mathbf{S}_{1}{-}\mathbf{S}_{2}\,|\,P\mathbf{R}_{1})\;D^{\mathbf{S}_{2}}_{\mu\nu}\right]$
@@
is the truncated Coulomb matrix with truncation radius $r_{\text{SC}}\hskip 1.69998pt{=}\hskip 1.69998pt\min_{j}N_{j}|\mathbf{a}_{j}|/2$, i.e., half of the shortest primitive translation vector of the supercell SC (29[Spencer2008, Irmler2018]. Note that we restrict all lattice vector differences appearing in Eq. (56) and (59) to the SC (29), i.e., $\mathbf{S}_{1}{-}\mathbf{S}_{2},\mathbf{R}_{1}{-}\mathbf{R}_{2}\hskip 1.69998pt{\in}\hskip 1.69998pt\text{SC}$.
@@
The primary difference between our algorithm and the $\Gamma$-point-only $GW$ implementation using Gaussian basis functions reported in Ref. [Graml2024] lies in the treatment of periodic boundary conditions. In our approach, both the irreducible density response function $\chi$ and the self-energy $\Sigma$ are computed via explicit lattice summations [Eqs. (42) and (56)] over all unit cells. In contrast, Ref. [Graml2024] restricts the lattice sum to the nearest-neighbor cell, effectively circumventing the need for a full lattice summation. This approach is only exact in the limit of large unit cells, and the presented $GW$ algorithm enables the treatment of small unit cells.